# Asymptotic and compound e-values: multiple testing and empirical Bayes

### v3

@@ 6.1.2 Optimal discovery compound e-values via empirical Bayes
where we recall that $\mathcal{G}$ is the class of all distributions supported on $[\underaccent{\bar}{\sigma}^{2},\bar{\sigma}^{2}]$. Moreover we let $\widehat{Q}=\widehat{H}\otimes\widehat{G}$ where $\widehat{H}$ is a distribution on $\mathbb{R}$ estimated based on $X_{1},\ldots,X_{K}$ and that is supported on $[\underaccent{\bar}{\lambda},\bar{\lambda}]\subseteq(0,\infty)$, where $\underaccent{\bar}{\lambda},\bar{\lambda}$ are fixed.
@@
6.1.3 Simulation study
We demonstrate the improved power of our two constructions through a simulation study. We simulate from the normal t-test model with $K=2000$ and let $n\in\{5,10\}$ be a simulation parameter. We consider two settings for $\sigma_{k}^{2}$, either all $\sigma_{k}^{2}=1$ or $\sigma_{k}^{2}\,{\buildrel\text{iid}\over{\sim}\,}\mathrm{Unif}[0.5,2]$. Moreover we set $\lambda_{k}=0$ for $1800$ hypotheses (nulls), and $\lambda_{k}=\xi/\sqrt{n}$ for the other 200 hypotheses. We call $\xi$ the effect size and vary it in the interval $[2,6]$ in our simulations.
  1. 1.
  2. 2.
  1. 3.
  2. 4.
  3. 5.
    Universal inference (UI): As in (11) but with numerator given by $\int p_{\lambda,\sigma^{2}}(x)\mathrm{d}(H\otimes\widehat{G})(\lambda,\sigma^{2})$.
  4. 6.
Figure 1: Power comparison of six multiple testing procedures under different sample sizes ($n\in\{5,10\}$) and variance settings ($\sigma_{k}^{2}=1$ vs. $\sigma_{k}^{2}\sim\mathrm{Unif}[0.5,2]$). Methods include Z-test (Oracle), Separable (Oracle), EB (Empirical Bayes), CUI (Compound Universal Inference), t-test, and UI (Universal Inference). Among data-driven methods (i.e., methods that do not have oracle knowledge regarding the variances), our proposed EB method has by far the largest power. Moreover, EB is indistinguishable from its oracle counterpart (EB-Oracle).
@@ 6.2.2 A prototypical construction of asymptotic compound e-values
7 An alternate definition of approximate compound p-values and e-values
In this section we provide an alternative definition in lieu of Definition 3.13 for approximate compound p-variables and e-variables. This definition has the advantage that it is easier to satisfy, but it provides weaker guarantees. We include it because the asymptotic notion it leads to for compound p-variables was central to the work of Ignatiadis and Sen (2025).
  1. (i)
  2. (ii)
When all $\mathbb{P}\in\mathcal{P}$ are atomless and $K=1$, then Definitions 7.1 and 3.13 are equivalent to each other and equivalent to Definition 3.1 on approximate p-variables and e-variables. The definitions also agree when $\delta=0$, for instance, $(\varepsilon,0)^{*}$-approximate compound e-variables are identical to $(\varepsilon,0)$-approximate compound e-variables.
In analogy to Definition 3.17, we define (uniformly) *asymptotic compound p-variables and e-variables, requiring, e.g., that $\lim_{m\to\infty}\varepsilon_{m}(\mathbb{P})=\lim_{m\to\infty}\delta_{m}(\mathbb{P})=0$ for each $\mathbb{P}\in\mathcal{P}$. We do not define strongly *asymptotic compound p-variables (or e-variables) since the natural definition is equivalent (under the atomless property) to strongly asymptotic compound p-variables (or e-variables).
Definition 7.1 is implied by Definition 3.13, as can be seen most clearly for approximate compound e-variables by comparing to the result of Proposition 3.15. Moreover, Proposition 3.14 has the following analogous result (whose proof we omit).
Proposition 7.2.
Proposition 7.3.
Theorem 4.4 demonstrated that approximate compound e-variables are naturally compatible with approximate FDR control when using the e-BH procedure. An analogous statement is not true for approximate compound e-variables without further conditions. However it is true in a regime where e-BH makes sufficiently many discoveries. The statement of the following proposition is inspired by Li and Barber (2019, Theorem 2), whose FDR bound also depends on the tail probability of the number of rejections.
Proposition 7.4.
@@ 8 Summary
We briefly point out that the term compound p-values has previously been used in a way that is distinct from Definition 2.1. Habiger and Peña (2014) considered a setting akin to the sequence model in Section 5 wherein $X=(X_{k}:k\in\mathcal{K})$ and in principle $X_{k}$ can be used to test $\mathcal{P}_{k}$ for all $k\in\mathcal{K}$. Let $P_{1},\dots,P_{K}$ be $[0,1]$-valued random variables. Habiger and Peña (2014) called these “compound p-values” if they are bona fide p-values (that is $\mathbb{P}(P_{k}\leq t)\leq t$ for all $t\in[0,1]$ and all $\mathbb{P}\in\mathcal{P}_{k}$) and depend also on data for other hypotheses $(X_{\ell}:\ell\neq k)$, i.e., $P_{k}=P_{k}(X)$ rather than $P_{k}=P_{k}(X_{k})$. In other words, “compound p-values” are not separable as per Definition 5.1, and so could be called “non-separable p-values” instead. The definition treats separable p-values as the default data analysis choice (which is reasonable given how p-values are often used in practice), and one needs a qualifier to refer to non-separable p-values. By contrast, we prefer to think of (compound) p-values and e-values as being non-separable by default (given their flexibility and generality), and to use an explicit qualifier for the separable case. Moreover, we use the term compound in relation to the averaging (compounding) operation inherent in Definition 2.1.
@@ Acknowledgments.
This work was completed in part with resources provided by the University of Chicago’s Research Computing Center.
Appendix A Compound e-values as weights in merging p-values
@@
Appendix B Proofs for Section 2
B.1 Proof of Theorem 2.3
@@ Appendix B Proofs for Section 2
Appendix C Proofs for Section 3
C.1 Proof of Proposition 3.2
@@ Appendix C Proofs for Section 3
Appendix D Proofs for Section 4
D.1 Proof of Theorem 4.5
Proof.
The first part directly follows from Theorems 4.2 and 4.4. To show the last statement, consider an FDR procedure $\mathcal{D}$ at level $\alpha$ and take the compound e-values $E_{1},\dots,E_{K}$ from (4). Define
@@ Appendix D Proofs for Section 4
Appendix E Proofs for Section 5
E.1 Proof of Proposition 5.4
@@ Appendix E Proofs for Section 5
$\sum_{k\in\mathcal{K}}\mathbb{E}^{\mathbb{P}_{k}}[s(X_{k})]=K.$
@@
Thus, if we define $q_{k}(x)=s(x)\overline{p}^{\circ}(x)$ for all $k\in\mathcal{K}$, we find that $q_{k}$ is indeed a $\mathrm{d}\nu$-density and $s(x)$ may be represented as in (10). ∎
E.2 Proof of Proposition 5.5
@@
Appendix F Proofs for Section 6
F.1 Proof of Theorem 6.3

### v4

@@ 6.1.2 Optimal discovery compound e-values via empirical Bayes
where we recall that $\mathcal{G}$ is the class of all distributions supported on $[\underaccent{\bar}{\sigma}^{2},\bar{\sigma}^{2}]$ and $f_{G}(\hat{\sigma}_{k}^{2})$ is defined in (19). Moreover we let $\widehat{Q}=\widehat{H}\otimes\widehat{G}$ where $\widehat{H}$ is a distribution on $\mathbb{R}$ estimated based on $X_{1},\ldots,X_{K}$ and that is supported on $[\underaccent{\bar}{\lambda},\bar{\lambda}]\subseteq\mathbb{R}$, where $\underaccent{\bar}{\lambda},\bar{\lambda}$ are fixed.
@@
We demonstrate the improved power of our two constructions through a simulation study in Section 7.
@@ 6.2.2 A prototypical construction of asymptotic compound e-values
7 Simulation study
The goal of this section is to empirically demonstrate that the approximate/asymptotic compound e-values of Section 6.1 can lead to a substantial power increase for the e-BH procedure. We follow the notation and setting of Section 6.1.
We simulate from the simultaneous normal t-test model with $K=2000$ and let $n\in\{5,10\}$ be a simulation parameter. We consider two settings for $\sigma_{k}^{2}$, either all $\sigma_{k}^{2}=1$ or $\sigma_{k}^{2}\,{\buildrel\text{iid}\over{\sim}\,}\mathrm{Unif}[0.5,2]$. Moreover we set $\lambda_{k}=0$ for $1800$ hypotheses (nulls), and $\lambda_{k}=\xi/\sqrt{n}$ for the other 200 hypotheses. We call $\xi$ the effect size and vary it in the interval $[2,6]$ in our simulations.
  1. 1.
  2. 2.
  1. 3.
  2. 4.
  3. 5.
    Universal inference (UI): As in (11) with numerator given by $\int p_{\lambda,\sigma^{2}}(x)\mathrm{d}(H\otimes\widehat{G})(\lambda,\sigma^{2})$.
  4. 6.
Figure 1: Power comparison of six multiple testing procedures under different sample sizes ($n\in\{5,10\}$) and variance settings ($\sigma_{k}^{2}=1$ vs. $\sigma_{k}^{2}\sim\mathrm{Unif}[0.5,2]$). Methods include z-Oracle, EB-Oracle, EB (Empirical Bayes), CUI (Compound Universal Inference), t-test, and UI (Universal Inference). Among data-driven methods (i.e., methods that do not have oracle knowledge regarding the variances), our proposed EB method has by far the largest power. Moreover, EB is indistinguishable from its oracle counterpart (EB-Oracle).
@@ Acknowledgments.
This work was completed in part with resources provided by the University of Chicago’s Research Computing Center. RW is supported by the Natural Sciences and Engineering Research Council of Canada (CRC-2022-00141, RGPIN-2024-03728). NI gratefully acknowledges support from the U.S. National Science Foundation (NSF DMS-2443410).
Appendix A Compound p-value definition in Habiger and Peña [2014]
We briefly point out that the term compound p-values has previously been used in a way that is distinct from Definition 2.1Habiger and Peña [2014] considered a setting akin to the sequence model in Section 5 wherein $X=(X_{k}:k\in\mathcal{K})$ and in principle $X_{k}$ can be used to test $\mathcal{P}_{k}$ for all $k\in\mathcal{K}$. Let $P_{1},\dots,P_{K}$ be $[0,1]$-valued random variables. Habiger and Peña [2014] called these “compound p-values” if they are bona fide p-values (that is $\mathbb{P}(P_{k}\leq t)\leq t$ for all $t\in[0,1]$ and all $\mathbb{P}\in\mathcal{P}_{k}$) and depend also on data for other hypotheses $(X_{\ell}:\ell\neq k)$, i.e., $P_{k}=P_{k}(X)$ rather than $P_{k}=P_{k}(X_{k})$. In other words, “compound p-values” are not separable as per Definition 5.1, and so could be called “non-separable p-values” instead. The definition treats separable p-values as the default data analysis choice (which is reasonable given how p-values are often used in practice), and one needs a qualifier to refer to non-separable p-values. By contrast, we prefer to think of (compound) p-values and e-values as being non-separable by default (given their flexibility and generality), and to use an explicit qualifier for the separable case. Moreover, we use the term compound in relation to the averaging (compounding) operation inherent in Definition 2.1.
Appendix B More on *approximate compound p-values and e-values
Here we elaborate further on Definition 3.17. In analogy to Definition 3.18, we define (uniformly) *asymptotic compound p-variables and e-variables, requiring, e.g., that $\lim_{m\to\infty}\varepsilon_{m}(\mathbb{P})=\lim_{m\to\infty}\delta_{m}(\mathbb{P})=0$ for each $\mathbb{P}\in\mathcal{P}$. We do not define strongly *asymptotic compound p-variables (or e-variables) since the natural definition is equivalent (under the atomless property) to strongly asymptotic compound p-variables (or e-variables).
Proposition 3.14 has the following analogous result (whose proof we omit).
Proposition B.1.
Proposition B.2.
Proof.
Hence, further arguing as in the proof of Theorem 2.3, for any $t>0$,
Theorem 4.3 demonstrated that approximate compound e-variables are naturally compatible with approximate FDR control when using the e-BH procedure. An analogous statement is not true for approximate compound e-variables without further conditions. However it is true in a regime where e-BH makes sufficiently many discoveries. The statement of the following proposition is inspired by Li and Barber [2019, Theorem 2], whose FDR bound also depends on the tail probability of the number of rejections.
Proposition B.3.
Proof.
Appendix C Compound e-values in multiple testing: addendum
C.1 Combination and derandomization
The connections between FDR controlling procedures, e-BH, and compound e-values laid out in Section 4 motivate the following general and practical mechanism for combining discoveries across multiple testing procedures. To be concrete, suppose we run $L$ different multiple testing procedures $\mathcal{D}_{1},\dots,\mathcal{D}_{L}$. The $\ell$-th procedure is applied to a subset of hypotheses $\{\mathcal{P}_{k}:\;k\in\mathcal{\mathcal{missing}}{S}_{\ell}\}$ where $\mathcal{S}_{\ell}\subseteq\mathcal{K}$ and controls the FDR at level $\alpha_{\ell}$. Then we may proceed as follows:
  1. 1.
    For the $\ell$-th multiple testing procedure, form compound e-values $E_{k}^{(\ell)},\,k\in\mathcal{S}_{\ell}$, for $\mathcal{P}_{k},\,k\in\mathcal{S}_{\ell}$ (which always exist by the proof of Theorem 4.4). They could be—but need not necessarily be—the implied ones formed in (4) . For $k\in\mathcal{K}\setminus\mathcal{S}_{\ell}$, set $E_{k}^{(\ell)}=1$.
  2. 2.
  3. 3.
The above construction is guaranteed to control the FDR at level $\alpha$, as long as all the individual procedures $\mathcal{D}_{\ell}$ form valid compound e-values. Note that the value of $\alpha_{\ell}$ does not matter, as it is only used in the construction of $E_{1}^{(\ell)},\dots,E_{K}^{(\ell)}$ , possibly implicitly (e.g., in the proof Theorem 4.4).
One important application of the above recipe is derandomization. Suppose that $\mathcal{D}_{\ell}$ is a randomized multiple testing procedure, that is, it is a function of both the data $X$ as well as a random variable $U_{\ell}$ generated during the analysis. Such randomness may not be desirable, since different random number generation seeds will lead to different sets of discoveries. In such cases, the above recipe can be used to construct a new derandomized procedure $\mathcal{D}$ that is less sensitive to $U_{1},\dots,U_{L}$ (formally, full derandomization would occur if $U_{\ell}$ are iid and $L\to\infty$). In this way Ren and Barber [2024] were able to derandomize the model-X knockoff filter [Candès et al., 2018], a flexible set of methods for variable selection in regression with finite-sample FDR control that previously relied on additional randomness.
A further application includes the following: Banerjee et al. [2023] used the above recipe for meta-analysis in which the $\ell$-th study only reports the set of tested hypotheses, the set of discoveries, as well as the targeted FDR level.
We end this subsection by using our recipe to derandomize the randomized e-BH procedures (including $R_{1}$-e-BH, $R_{2}$-e-BH, $R$-e-BH, and $U$-e-BH) proposed by Xu and Ramdas [2024]. For instance, $U$-e-BH draws $U\sim\mathrm{Unif}[0,1]$ and then applies e-BH to $E_{1}/U,\ldots,E_{K}/U$. Denote any of these procedures by $\mathcal{D}$. All of these procedures have in common the following properties: they operate on e-values $E_{1},\ldots,E_{K}$ as well as an external source of randomization $U$, they control the FDR without requiring further additional assumptions over e-BH, and finally they make at least as many discoveries as e-BH (that is, $\mathcal{D}\supseteq\mathcal{D}_{\text{e-BH}}$ almost surely), often making strictly more discoveries. Both Xu and Ramdas [2024] and Lee and Ren [2024] have noted the challenge of derandomizing these procedures; naive attempts lead back to e-BH applied to the original e-values. Now consider applying the derandomization procedure of this section with $\mathcal{D}_{\ell}$ denoting an application of $\mathcal{D}$ with external randomization seed $U_{\ell}$, $S_{\ell}=\mathcal{K}$, and weights $w_{\ell}=1$. Then the resulting procedure controls the FDR and is different from e-BH.
C.2 Compound e-values as weights in p-value based multiple testing
Ignatiadis et al. [2024] explained that e-values can be used as weights in p-value based multiple testing. We explain why compound e-values are the natural notion of weights for multiple testing with the p-BH procedure. We first define the p-BH procedure of Benjamini and Hochberg [1995] for FDR control based on p-values, as well as weighted generalizations thereof.
Definition C.1 (p-BH, weighted p-BH, ep-BH).
with the convention $\max(\varnothing)=0$. When $W_{k}=1$ for all $k$, then the above procedure is the p-BH procedure of Benjamini and Hochberg [1995]. When $W_{k}=w_{k}$ is deterministic and such that $\sum_{k\in\mathcal{K}}w_{k}=K$, then the above procedure is the weighted p-BH procedure of Genovese et al. [2006]. When $W_{k}=E_{k}$ and $E_{1},\dots,E_{K}$ are compound e-variables, then the above procedure is the ep-BH (e-weighted p-BH) procedure of Ignatiadis et al. [2024].
The generalization of p-BH to allow for deterministic weights by Genovese et al. [2006] was motivated by differentially prioritizing hypotheses to improve the power of p-BH while maintaining frequentist FDR control. The requirement $\sum_{k\in\mathcal{K}}w_{k}=K$ may be interpreted as a fixed size budget to be split across hypotheses. When the p-values are positive regression dependent on a subset (PRDS, Benjamini and Yekutieli [2001]) under the distributions in $\mathcal{P}$ and the weights are deterministic then existing arguments in the literature [Ramdas et al., 2019] demonstrate that the FDR of the weighted p-BH procedure is controlled at $(\alpha/K)\sum_{k\in\mathcal{N}}w_{k}$ under this assumption, which is bounded by $\alpha$ under the budget constraint $\sum_{k\in\mathcal{K}}w_{k}=K$. If the weights $W_{1},\dots,W_{K}$ are random and independent of $P_{1},\dots,P_{K}$, then by applying the above argument conditional on the weights and by iterated expectation, it follows that the FDR is controlled at $\sup_{\mathbb{P}\in\mathcal{P}}(\alpha/K)\mathbb{\sum}_{k:\mathbb{P}\in\mathcal{P}_{k}}\mathbb{E}^{\mathbb{P}}[W_{k}]$. The latter is controlled at $\alpha$ precisely when $W_{1},\dots,W_{K}$ are compound e-values. We note that Ignatiadis et al. [2024] defined the ep-BH procedure and proved FDR control (in their Theorem 4) when $E_{1},\dots,E_{K}$ are e-values. However, Ignatiadis et al. [2024, Section 4.5] explained that the guarantees would continue to hold for compound e-values (although their definition is slightly different from ours).
C.3 Compound e-values as weights in merging p-values
@@
C.4 Compound p-values in multiple testing
So far in this section we have focused on the role of compound e-values in multiple testing. Here we briefly discuss some known results about compound p-values. Armstrong [2022] provided a counterexample of jointly independent compound p-values $P_{1},\dots,P_{K}$, such that the p-BH procedure (Definition C.1) does not control the FDR. On the other hand, we show the following result, slightly generalizing a claim by Armstrong [2022].
Proposition C.3.
If $P_{1},\dots,P_{K}$ are (asymptotic) compound p-values, then the p-BY [Benjamini and Yekutieli, 2001] procedure (that is, the p-BH procedure applied at level $\alpha^{\prime}=\alpha/\sum_{k\in\mathcal{K}}1/k$) provides (asymptotic) control of the FDR at $\alpha$ under arbitrary dependence among $P_{1},\dots,P_{K}$.
Proof.
The nonasymptotic result was proved by Armstrong [2022], but here is an alternate proof that makes it transparent to extend to the asymptotic case. Wang and Ramdas [2022] proved that the p-BY procedure may be equivalently described by two steps. First, the p-values $P_{k}$ are calibrated to e-values $E_{k}$ via a specific choice of calibrator $f:p\mapsto T(\alpha/(\ell_{K}p))/\alpha$, where
Second, the e-BH procedure is applied to $E_{1},\dots,E_{K}$. With this, the nonasymptotic claim then follows directly from Theorems 2.3 and 4.3. The asymptotic claim follows because calibrating asymptotic p-values yields asymptotic e-values, which when used with the e-BH procedure provides asymptotic FDR control. ∎
Furthermore, Armstrong [2022] proved that the p-BH procedure with compound p-values controls the FDR under weak dependence asymptotics (as in Storey et al. [2004]); such a property of the p-BH procedure (with compound p-values) was also anticipated earlier by Efron [2007].
Appendix D Optimal simple separable compound e-values for general utility functions
Here we present a result akin to Theorem 5.2 for more general utility functions $U(\cdot)$ going beyond $U(\cdot)=\log(s)$. Our notation follows that of Section 5.2. Fix a utility function $U:[0,\infty]\to[-\infty,\infty]$. We seek to solve the following optimization problem.
$\underset{s(\cdot)}{\text{maximize}}$ $\frac{1}{K}\sum_{k\in\mathcal{K}}\mathbb{E}^{\mathbb{Q}_{k}^{\circ}}[U(s(X_{k}))]$ (28)
subject to $s:\mathcal{X}\to[0,\infty]$
[…]
Note that optimization problem (9) is a special case of optimization problem (28) for the choice $U(\cdot)=\log(\cdot)$. The proof of the next theorem is analogous to that of Theorem 5.2 and thus omitted.
Theorem D.1 (Optimal simple separable compound e-values; general utility).
Let ¯$\overline{\mathbb{P}}^{\circ}=\sum_{k\in\mathcal{K}}\mathbb{P}_{k}^{\circ}/K$ and ¯$\overline{\mathbb{Q}}^{\circ}=\sum_{k\in\mathcal{K}}\mathbb{Q}_{k}^{\circ}/K$. Suppose that $s^{*}(\cdot)$ solves the following optimization problem:
$\underset{s(\cdot)}{\textnormal{maximize}}$ $\mathbb{E}^{X^{\prime}\sim\overline{\mathbb{Q}}^{\circ}}[\log(s(X^{\prime}))]$ (29)
subject to $s:\mathcal{X}\to[0,\infty]$
Then $s^{*}(\cdot)$ also solves optimization problem (28) and thus yields the optimal simple separable compound e-values for the utility function $U(\cdot)$.
By Theorem D.1 we can use existing results for problem (29) to solve problem (28). Below we apply our theorem to solutions from Koning [2025], who provides optimal functions for problem (29) under several choices of $U(\cdot)$. It will be convenient to introduce the notation,
$s^{\mathrm{ODP}}(x)=\frac{\sum_{j\in\mathcal{K}}q_{j}^{\circ}(x)}{\sum_{j\in\mathcal{K}}p_{j}^{\circ}(x)}$
for the optimal discovery compound e-value function in (10).
  • Generalized means. Consider $U(s)=U_{h}(s)=(s^{h}-1)/h$ for $h\in(0,1)$. [^10] Note that as $h\to 0$, $U_{h}(s)\to\log(s)$, i.e., we recover the logarithmic utility function. Moreover suppose that $\mathbb{E}^{X^{\prime}\sim\overline{\mathbb{P}}^{\circ}}[(s^{\mathrm{ODP}}(x))^{1/(1-h)}]\in(0,\infty)$. Then a solution to (28) is given by:
    $s^{*}(x)=s^{*}_{h}(x)=\frac{\left(s^{\mathrm{ODP}}(x)\right)^{1/(1-h)}}{\mathbb{E}^{X^{\prime}\sim\overline{\mathbb{P}}^{\circ}}\left[\left(s^{\mathrm{ODP}}(x)\right)^{1/(1-h)}\right]}.$
  • Generalized means with clipping. Fix $M\in(0,\infty)$ and $h\in(0,1)$ and define the utility $U(s)=U_{h,M}(s)=U_{h}(s\wedge M)$ with $U_{h}(\cdot)$ defined in the above bullet point. This utility function clips the argument at $M$, so that values exceeding $M$ provide no additional utility beyond what $M$ provides. Then, provided that $\mathbb{P}_{k}(s(X_{k})>0)>0$ for at least one $k\in\mathcal{K}$, a solution to (28) is given by:
    $s^{*}(x)=s^{*}_{h,M}(x)=\frac{\left(s^{\mathrm{ODP}}(x)\right)^{1/(1-h)}\wedge M}{\mathbb{E}^{X^{\prime}\sim\overline{\mathbb{P}}^{\circ}}\left[\left(s^{\mathrm{ODP}}(x)\right)^{1/(1-h)}\wedge M\right]}.$
Appendix E Proofs for Section 2
E.1 Proof of Theorem 2.3
@@ Appendix E Proofs for Section 2
Appendix F Proofs for Section 3
F.1 Proof of Proposition 3.2
@@ Appendix F Proofs for Section 3
Let us explicitly define the two statements under consideration:
  1. (E1)
    It holds that
  2. (E2)
We seek to show the equivalence of (E1) and (E2).
@@
Appendix G Proofs for Section 4
G.1 Proof of Theorem 4.4
Proof (continued)..
To show the last statement, consider an FDR procedure $\mathcal{D}$ at level $\alpha$ and take the compound e-values $E_{1},\dots,E_{K}$ from (4). Define
@@ Appendix G Proofs for Section 4
Appendix H Proofs for Section 5
H.1 Proof of Proposition 5.4
@@ Appendix H Proofs for Section 5
$\sum_{k\in\mathcal{K}}\mathbb{E}^{\mathbb{P}_{k}^{\circ}}[s(X_{k})]=K.$
@@
Thus, if we define $q_{k}^{\circ}(x)=s(x)\overline{p}^{\circ}(x)$ for all $k\in\mathcal{K}$, we find that $q_{k}^{\circ}$ is indeed a $\mathrm{d}\nu$-density and $s(x)$ may be represented as in (10). ∎
H.2 Proof of Proposition 5.5
@@
Appendix I Proofs for Section 6
I.1 Proof of Theorem 6.3